You invested a total of $25,000 at 3% and 4 1/2% simple interest. During one year, the two accounts earned $900. How much did you invest in each account?
"thank you so much. i really want to understand this investments."

Answer by richwmiller(17219)

x+y=25000,
0.03*x+0.045*y=900
x=25000-y
0.03*(25000-y)+0.045*y=900
750-0.03y+0.045*y=900
0.015*y=150
y= 10000
x=25000-y
x=15000
check
0.03*15000+0.045*10000=900
450+450.0=900
900.0=900
ok